Grade 11Physical Chemistry

Find molality and molarity of 15% solution of sulphuric acid when it's density is 1.10g/cm^3 and molar mass= 98amu

Sparsh
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Sparsh
 
15% H₂SO₄ solution means 15 g of H₂SO₄ is present in 100 g of Water.
No. of moles of H₂SO₄ = 15 g / 98 g/mol = 0.153 mol
Weight of solvent = 100g - 15g = 85g = 85 × 10^-3 kg
 
Molality = Moles of solute / Weight of Solvent in kg
= 0.153 / (85 × 10^-3)
= 1.8 m
Molality of given solution is 1.8 m ……[1]
 
Volume of solution = Mass of solution / Density
= 100 g / 1.1 g/cm^3
= 90.90 cm^3
= 0.090 Litres
 
Molarity = Moles of solute / Volume of solution in litres
= 0.153 / 0.090
≈ 1.7 M
Molarity of given solution is 1.7 M 
 
Regards
Arun (askIITians forum expert)